OpenStudy (anonymous):

Help w/ math question??? "Let g(x)=x^2 + 2x - 6. Find the value of 2[g(2)]" So far I've been able to solve all the other ones similar to it, but this one has me stuck for some reason?? Could someone please explain how to solve and whatnot? Thanksxx

OpenStudy (anonymous):

g(2)=2^2+2*2-6=2. 2g(2) = 2*2 Does that make sense? First compute g(2). Then multiply that by 2 to get 2g(2)

OpenStudy (anonymous):

Yeah, I think I get it now. Thanks!